OpenStudy (anonymous):

Hi I'm learning the Laws of Sine and I don't understand part of it. 7.7/sin 43.2 degrees = 9.1/sin B → sin B = 9.1 (sin 43.2 degrees)/7.7 = 0.8090 → B = 54° or B = 180° - 54° = 126° (to the nearest tenth). How do they go from 0.8090 to B=54.

OpenStudy (anonymous):

I feel like it's simple but I can't remember.

OpenStudy (luffingsails):

Law of Sines: a/sin A = b/sin B = c/sin C Right?

OpenStudy (anonymous):

yes

OpenStudy (luffingsails):

Okay, just checking. lol Can you use a calculator?

OpenStudy (luffingsails):

for the problem, I mean.

OpenStudy (anonymous):

Yea I can

OpenStudy (anonymous):

Well that's just it I'm not sure what I'm supposed to be calculating

OpenStudy (luffingsails):

Okay, so they finish with sin B = 9.1 (sin 43.2 degrees)/7.7 = 0.8090 Take the inverse sine of .8090

OpenStudy (anonymous):

Oh that makes so much more sense thanks I get it now.

OpenStudy (luffingsails):

Sure. glad to help.